OpenStudy (anonymous):

((8x+13)+(x2-5x))(x) how would you solve for that???

OpenStudy (abb0t):

Distribute the "x" first.

OpenStudy (anonymous):

I thought whatever is in the parentheses, is like solved first?

OpenStudy (mathstudent55):

Is this the multiplication of the quantity and x? \( \large [(8x+13)+(x^2-5x)]x\)

OpenStudy (anonymous):

i do not think so

OpenStudy (mathstudent55):

Then what is the problem?

OpenStudy (anonymous):

Btw, its problem number 2

OpenStudy (mathstudent55):

I got it.

OpenStudy (mathstudent55):

Here are the two functions f and g. \(f(x) = 8x + 13\) \(g(x) = x^2 - 5x\) 2. \((f + g)(x) = f(x) + g(x)\) \((f + g)(x) = 8x + 13 + x^2 - 5x\) \((f + g)(x)= x^2 + 3x +13\)

OpenStudy (mathstudent55):

All you need to do for 2. is to add the functions f and g.

OpenStudy (anonymous):

ooooh okay, THANKYOU SO MUCH!!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

what about number 11? How would you find the composite function? do you plug in and solve or something else?

OpenStudy (mathstudent55):

For composite functions, plug in the inner function into x of the outer function. 11. \(f(g(x)) = f(2x - 3) = (2x - 3)^2\)